# Revealing the $\chi_{\rm eff}$-$q$ correlation among Coalescing Binary Black Holes and Evidence for AGN-driven Hierarchical Mergers

### v1

@@ Abstract
The origin of the correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of LIGO-Virgo-KAGRA’s binary black holes (BBHs) is still an open question. Motivated by recent identification of two subpopulations of the BBHs, in this work we investigate the potential $\chi_{\rm eff}-q$ correlation for each subpopulation. Surprisingly, the $\chi_{\rm eff}$-$q$ correlation vanishes for the low-mass subpopulation if we introduce a second $\chi_{\rm eff}$ distribution for the high-mass subpopulation likely originating from hierarchical mergers. The first subpopulation has a narrow $\chi_{\rm eff}$ distribution peaking at $\sim 0.05$, whose primary-mass function cuts off at $\sim 45M_{\odot}$, in agreement with first-generation BBHs. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\chi,2}=0.35^{+0.18}_{-0.22}$, consistent with the expectation of hierarchical mergers formed in the disks of active galactic nucleus (AGNs). We infer $\mu_{\chi,2}>0$ at 98.7% credible level, and a symmetric $\chi_{\rm eff}$ distribution for the second subpopulation is disfavored by $\mathcal{B}\sim 5$. However, negative values of $\chi_{\rm eff}$ are also measured, indicating that the hierarchical mergers may take place within both star clusters and AGN disks. We find a Bayes factor of $\ln\mathcal{B}=5.2$ for two distinct $\chi_{\rm eff}$ distributions relative to single $\chi_{\rm eff}$ distribution that conditioned on mass ratios. Therefore we conclude that the $\chi_{\rm eff}$-$q$ correlation in the entire population can be explained as the superposition of two subpopulations. Additionally, we suggest to use a flexible mass function to reduce the bias in $\chi_{\rm eff}$-$q$ correlation that may be introduced by model mis-specification.
@@ 1 Introduction
Callister et al. (2021), for the first time, reported the anti-correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of BBHs with data from GWTC-2 (Abbott et al., 2019, 2021a). This anti-correlation was subsequently confirmed by Abbott et al. (2023a) with data from GWTC-3 (Abbott et al., 2019, 2021a, 2024, 2023b). More detailed analysis (e.g., Adamcewicz & Thrane, 2022; Adamcewicz et al., 2023) had shown that the $\chi_{\rm eff}$-$q$ anti-correlation is not artificial. However, the origin of this anti-correlation is still in debate (Callister et al., 2021; Abbott et al., 2023a).
Many simulations suggested that hierarchical mergers in the disks of active galactic nucleus (AGNs) could explain the observed $\chi_{\rm eff}$-$q$ correlation (e.g., McKernan et al., 2022; Santini et al., 2023; Cook et al., 2024). The spin orientations and orbital angular momenta of binary black holes (BBHs) will be modulated by the disks, causing the mergers to favor positive effective spins. Additionally, the migration traps in AGN disks will produce mergers involving multiple generations of black holes (BHs) that have unequal masses and larger (positive) effective spins. Previously, we found that the coalescing BHs can be divided into two subpopulations with significantly different spin-magnitude versus component-mass distributions, which are nicely consistent with first- and higher-generation BHs (Li et al., 2024b). We also found that a fraction of the the hierarchical mergers have aligned spin tilts and asymmetric mass ratios, which may give rise to the $\chi_{\rm eff}$-$q$ anti-correlation.
@@
Therefore, in this study, we delve into the question of whether the anti-correlation between $\chi_{\rm eff}$-$q$ in BBHs originates from a superposition of various formation channels / subpopulations (Li et al., 2024b), if it emerges from the evolutionary processes of a single population with a common formation channel (e.g., Banerjee & Olejak, 2024), or alternatively, both effects have contributed to the observed correlation. Additionally, we investigate the $\chi_{\rm eff}$ distribution in the high-mass range that corresponds to the hierarchical mergers, in order to find out whether these events originate from the AGN disks, or alternatively from the star clusters which exhibits a symmetric $\chi_{\rm eff}$ distribution (Payne et al., 2024; Antonini et al., 2024).
This letter is organized as follows: In Section 2 and Section 3, we introduce the methods and the results. In Section 4, we present the conclusions and engage in discussions.
@@ 2 Methods
We first use the same $\chi_{\rm eff}-q$ distribution model as that used in Callister et al. (2021) to fit data of GWTC-3 for comparison. Specifically, the mean and width of the $\chi_{\rm eff}$ distribution change linearly with $q$ (referred to as the Base model); see Appendix B.2 for the detailed formula. The rate evolution model is the MD model (Madau & Dickinson, 2014), as defined in the Appendix B.3. To reduce the bias that may be introduced by the mis-modeling of primary-mass distribution, we use the concise non-/semi-parametric formula PowerLaw+Spline (PS; Edelman et al., 2022; Abbott et al., 2023a) instead of the popular parametric formula PowerLaw+Peak (PP; Talbot & Thrane, 2018; Abbott et al., 2021b), see Appendix B.1 for the details of mass functions. We find the PP model is disfavored compared to the PS model by $\ln\mathcal{B}=6$, due to its failure in modeling the additional structures in the primary-mass distribution (Li et al., 2021; Tiwari, 2022; Edelman et al., 2023; Abbott et al., 2023a; Callister & Farr, 2024). With the PP model, we find an anti-correlation between $\chi_{\rm eff}$ and $q$ distribution (i.e., $a<0$) at 99.2% credible level. However, the credibility decreases to 89.8% with the more flexible PS model, indicating that mis-modeling in the primary-mass distribution may introduce bias in measuring the $\chi_{\rm eff}$-$q$ correlation; however, the overall tendency remains unchanged, as shown in Figure 1.
@@
$\pi_{1}$ $(m_{1},\chi_{\rm eff}|{\bf\Lambda})=\mathcal{PS}(m_{1}|{\bf\Lambda})P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ (2)
@@
where $r_{2}$ is the mixed fraction of the second subpopulation, $\mathcal{PS}$ and $\mathcal{PL}$ are the PS and power-law primary-mass function, and $\mathcal{G}_{[-1,1]}$ is the truncated Gaussian distribution. $P(m_{2}|m_{1};{\bf\Lambda})$ is the normalized secondary-mass distribution conditioned on $m_{1}$, $P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ is the Base model that encodes the $\chi_{\rm eff}-q$ correlation (Callister et al., 2021), see Appendix B for the definition.
@@
$\pi_{\rm tran}$ $(m_{1},m_{2},\chi_{\rm eff}|{\bf\Lambda})=\mathcal{PS}(m_{1}|{\bf\Lambda})P(m_{2}|m_{1};{\bf\Lambda})$ (4)
@@ 3 Results
Base(PS) with $a=0$,$b=0$ 0.8
Transition(PS) 3.1
Transition(PS) with $a=0$,$b=0$ 5.2
Mixture(PS) 2.1
Mixture(PS) with $a=0$,$b=0$ 4
@@
Base(PP) with $a=0$,$b=0$6.9
Transition(PP) with $a=0$,$b=0$3
Note: All the log Bayes factors are relative to the Base model
that encode the $\chi_{\rm eff}$-$q$ correlation and accompanied with PS
@@
Figure 1: Constraints on the mean $\mu_{\chi}(q)$ and standard deviation $\sigma_{\chi}(q)$ of the $\chi_{\rm eff}$ distribution, as a function of BBH mass ratio $q$. The solid curves are the medians and the colored bands are the 90% credible intervals.
Figure 2: Posteriors of the hyperparameters describing the $\chi_{\rm eff}-q$ correlation. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
In Table 1, we summarize the Bayes factors of the novel models in this work compared to the Base model. The Transition model and Mixture model are more favored than the Base model, with $\ln\mathcal{B}=3.1$ and $\ln\mathcal{B}=2.1$. Furthermore, the Transition model and Mixture model without the $\chi_{\rm eff}$-$q$ correlation (i.e., $a=0,b=0$) are even more favored. Figure 2 shows the posteriors of the hyperparameters describing the $\chi_{\rm eff}$ and $q$ correlation, the parameters of the Transition model and Mixture model highly consistent with each other. In both models, the slope parameters $a$ and $b$ are consistent with zero, indicating there is no evidence for a $\chi_{\rm eff}-q$ correlation in the first subpopulation or the low-mass range. In subsequent analysis, we adopt the results inferred with $a=0,b=0$, although the posterior distributions of parameters are broadly consistent with the results inferred with variable $a$ and $b$, see Appendix C.
@@
We find that the second subpopulation shows a significantly different $\chi_{\rm eff}$ distribution compared to the first subpopulation, see Figure 3. The divergence point (of the Transition model) in the primary-mass function is $m_{\rm t}=49^{+13}_{-10}M_{\odot}$, with a transition scale of $6^{+4}_{-4}M_{\odot}$. As for the Mixture model, the maximum mass of the first subpopulation is $\sim 45M_{\odot}$, which is consistent with the lower edge of the Pair-instability Supernova BH mass gap (Farmer et al., 2019). The minimum mass of the second subpopulation is $\sim 30M_{\odot}$. These results are consistent with our previous analysis using spin magnitudes of BBHs (Wang et al., 2022; Li et al., 2024b), see Appendix C for the distributions of other related parameters. The $\chi_{\rm eff}$ distributions of both subpopulations are not symmetric with respect to zero, indicating the presence of formation channels other than star clusters.
@@
We find that $\mu_{\chi,2}>0$ at 98.7% credible level, and a symmetric $\chi_{\rm eff}$ distribution with respect to zero for the second population (i.e.,$\mu_{\chi,2}=0$) is disfavored by a Bayes factor of $\ln\mathcal{B}=1.6$ ($\mathcal{B}\sim 5$), which indicates that the hierarchical mergers cannot be produced (solely) by star clusters. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\rm eff,2}=0.34^{+0.17}_{-0.22}$, which is typical for the hierarchical mergers in the AGN disks (Yang et al., 2019). We have also inferred with variable edges ($\chi_{\rm min}$, $\chi_{\rm max}$) truncated on the second $\chi_{\rm eff}$ distribution, see Appendix D. Though the case $\chi_{\rm min}=0$ cannot be ruled out, the location of the lower edge prefers a negative value. Therefore, it is likely that the star clusters also contribute to the hierarchical mergers.
@@
Figure 3: Effective-spin distributions of the two subpopulation inferred by the Transition and Mixture models with $a=0$,$b=0$. The lines represent the mean values and 90% credible intervals
@@ 4 Conclusions and Discussion
In this work, we investigate the origins of $\chi_{\rm eff}$-$q$ correlation in the BBHs with data of GWTC-3. Different from the previous analysis (Callister et al., 2021; Abbott et al., 2023a), we adopt a more flexible primary-mass function (e.g., PS; Edelman et al., 2022). We find the evidence for the $\chi_{\rm eff}$-$q$ anti-correlation inferred with the PS model is weaker than that inferred with the PP model (Talbot & Thrane, 2018; Abbott et al., 2021b). This indicates that the parametric formula PP may introduce bias in the $\chi_{\rm eff}$-$q$ distribution, since there are additional structures in the primary-mass distribution beyond one power law plus one peak (Abbott et al., 2023a; Tiwari, 2022; Edelman et al., 2023; Callister & Farr, 2024).
With the dedicated models that introduce a second $\chi_{\rm eff}$ distribution for the second (high-spin) subpopulation, the inferred $\chi_{\rm eff}$-$q$ correlation vanishes. Therefore, we can conclude that the $\chi_{\rm eff}$-$q$ anti-correlation of the entire BBH population mainly results from the superposition of two subpopulations. Additionally, there is no evidence for a $\chi_{\rm eff}-q$ correlation in each subpopulation identified by the previous work (Li et al., 2024b, a), see Appendix E.
What are the second subpopulation?. We point to the hierarchical mergers in gas-rich environments, such as AGN disks (Yang et al., 2019). Our previous works (Wang et al., 2022; Li et al., 2024b) identified a subpopulation of hierarchical mergers in BBHs, providing smoking-gun evidence, specifically a spin-magnitude distribution of approximately 0.7 (Gerosa & Fishbach, 2021). Antonini et al. (2024) attributed this subpopulation to hierarchical mergers in star clusters, as they found that the $\chi_{\rm eff}$ distribution is consistent with a uniform distribution in the range of approximately (-0.5, 0.5). In this work, we employ a more flexible model and find that the high-spin subpopulation prefers an asymmetric $\chi_{\rm eff}$ distribution (with respect to zero) over a symmetric one, with a logarithmic Bayes factor of $\ln\mathcal{B}=1.6$ (where $\mathcal{B}\sim 5$). we also infer that the peak of the second $\chi_{\rm eff}$ distribution $\mu_{\chi,2}>0$ at 98.7% credible level.
Many simulations (e.g., Yang et al., 2019; Santini et al., 2023; Cook et al., 2024) show that the hierarchical mergers in disk-like environments can produce a $\chi_{\rm eff}$ distribution peaking at $\sim 0.35$, which is consistent with our results, $\mu_{\chi,2}=0.35^{+0.17}_{-0.22}(0.30^{+0.19}_{-0.19})$, from the Transition (Mixture) model. However, contribution of formation channels that can produce significantly negative $\chi_{\rm eff}$ values, such as hierarchical mergers in star clusters, cannot be ruled out, see Appendix D.
Kick velocities of the merger remnants are also helpful in determining the formation environments of hierarchical mergers (Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2024). Several population analyses suggested that the excess of BBHs in $\sim 30-40M_{\odot}$ are consistent with the dynamical formation channels (Wang et al., 2022; Li et al., 2024a; Ray et al., 2024), and the ratio of merger rates between hierarchical mergers and first-generation (dynamical) BBHs are consistent with the star cluster origins (Antonini et al., 2023). Therefore, it is possible that the hierarchical mergers detected by LIGO/Virgo/KAGRA originate from both formation environments. However, there are only $\sim 11$ hierarchical mergers (high-spin events) in GWTC-3 (Li et al., 2024b), making it difficult to determine the mixture fractions of AGN-disk like and star-cluster formation channels. We will address such issue when the GW data are significantly enriched.
@@
Does any single formation channel contributes to $\chi_{\rm eff}-q$ correlation? (Banerjee & Olejak, 2024) We find there is no evidence for a $\chi_{\rm eff}-q$ correlation in the low-mass sub-populations, which are associated with the first-generation BBHs (Li et al., 2024b). However, this sub-population may not originate from a single formation channel. Both isolated and dynamical formation channels have contributed to first-generation BBHs (Wang et al., 2022; Li et al., 2024b; Godfrey et al., 2023; Li et al., 2024a; Ray et al., 2024). Therefore, we further investigate the $\chi_{\rm eff}-q$ correlation of BBHs originating from potential formation channels (Wang et al., 2022; Li et al., 2024b; Godfrey et al., 2023; Li et al., 2024a; Ray et al., 2024) and find no evidence for either subpopulation. See Appendix E for details. Interestingly, we find a stronger $\chi_{\rm eff}-q$ correlation in the potential subpopulation associated with the dynamically formed BBHs (i.e., excluding the BBHs from isolated channels as indicated by Li et al. (2024a)). This indicates that the $\chi_{\rm eff}-q$ correlation originates from dynamical formation channels, and that the AGN-disk formation channels may take a significant contribution (Santini et al., 2023; Cook et al., 2024). The stable mass transfer formation channel is expected to produce BBHs with $\chi_{\rm eff}-q$ anti-correlation (Olejak et al., 2024; Banerjee & Olejak, 2024), while BBHs from the common envelop channel may exhibit an opposite correlation (Bavera et al., 2020, 2021). The $\chi_{\rm eff}-q$ correlation in the isolated channels may be measured when the detections are enriched and the sup-populations are determined.
@@
The fourth observing run (O4) of the LIGO-Virgo-KAGRA GW detectors is currently underway, and the number of detections is rapidly increasing (see https://gracedb.ligo.org/latest/). At the end of O4, more than four times as many events are expected to be observed compared to O3 (Callister, 2024). With the enriched data more subpopulations / formation channels of BBHs may be identified (Zevin et al., 2021), and the mixture fractions and the parameter correlations of subpopulations can be better determined (Li et al., 2022, 2024a; Guo et al., 2024; Heinzel et al., 2024; Callister, 2024).

### v2

# Revealing the $\chi_{\rm eff}$-$q$ correlation among Coalescing Binary Black Holes and Tentative Evidence for AGN-driven Hierarchical Mergers

@@ Abstract
The origin of the correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of LIGO-Virgo-KAGRA’s binary black holes (BBHs) is still an open question. Motivated by the recent identification of two subpopulations of the BBHs, in this work we investigate the potential $\chi_{\rm eff}-q$ correlation for each subpopulation. Surprisingly, the $\chi_{\rm eff}$-$q$ correlation either significantly weakens or disappears for the low-mass subpopulation if we introduce a second $\chi_{\rm eff}$ distribution for the high-mass subpopulation, which likely originates from hierarchical mergers. This suggests that the $\chi_{\rm eff}$-$q$ correlation in the overall population can be explained by the superposition of two distinct subpopulations. We find Bayesian evidence strongly favoring two separate $\chi_{\rm eff}$ distributions over a single mass-ratio-dependent distribution, with Bayes factors $\ln\mathcal{B}>4.2$. The first subpopulation has a narrow $\chi_{\rm eff}$ distribution peaking at $\sim 0.05$, whose primary-mass function showing a rapid decline beyond $\sim 40M_{\odot}$, in agreement with first-generation BBHs. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\chi,2}\sim 0.4$, aligning with predictions for hierarchical mergers in active galactic nucleus (AGN) disks. However, we cannot exclude negative $\chi_{\rm eff}$values in the second subpopulation, suggesting hierarchical mergers might occur both in AGN disks and stellar clusters. Furthermore, the inferred second $\chi_{\rm eff}$ distribution might alternatively arise from other formation channels, such as stable mass transfer or chemically homogeneous evolution, if not interpreted as hierarchical mergers.
@@ 1 Introduction
Callister et al. (2021), for the first time, reported the anti-correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of BBHs with data from GWTC-2 (Abbott et al., 2019, 2021a). This anti-correlation was subsequently confirmed by Abbott et al. (2023a); Adamcewicz et al. (2023) with data from GWTC-3 (Abbott et al., 2019, 2021a, 2024, 2023b), though the evidence is reduced with a more flexible model (Heinzel et al., 2024). However, the origin of this anti-correlation is still in debate (Callister et al., 2021; Abbott et al., 2023a).
Many simulations suggested that hierarchical mergers in the disks of active galactic nucleus (AGNs) could explain the observed $\chi_{\rm eff}$-$q$ correlation (e.g., McKernan et al., 2022; Santini et al., 2023; Cook et al., 2024). The spin orientations and orbital angular momenta of binary black holes (BBHs) will be modulated by the disks, causing the mergers to favor positive effective spins. Additionally, the migration traps in AGN disks will produce mergers involving multiple generations of black holes (BHs) that have unequal masses and larger (positive) effective spins. Previously, we found that the coalescing BHs can be divided into two subpopulations with significantly different spin-magnitude versus component-mass distributions, which are nicely consistent with first- and higher-generation BHs (Li et al., 2024b). We also found that a fraction of the population that we interpreted as hierarchical mergers have aligned spin tilts and asymmetric mass ratios, which may give rise to the $\chi_{\rm eff}$-$q$ anti-correlation.
@@
In this study, we delve into the question of whether the anti-correlation between $\chi_{\rm eff}$-$q$ in BBHs originates from a superposition of various formation channels / subpopulations (Li et al., 2024b), if it emerges from the evolutionary processes of a single population with a common formation channel (e.g., Banerjee & Olejak, 2024), or alternatively, both effects have contributed to the observed correlation. Additionally, we investigate the $\chi_{\rm eff}$ distribution in the high-mass range that might corresponds to the hierarchical mergers, in order to find out whether these events originate from the AGN disks, or alternatively from the star clusters which exhibits a symmetric $\chi_{\rm eff}$ distribution (Payne et al., 2024; Antonini et al., 2025).
This work is organized as follows: In Section 2 and Section 3, we introduce the methods and the results. In Section 4, we present the conclusions and engage in discussions.
@@ 2 Methods
We first use the same $\chi_{\rm eff}-q$ distribution model as that used in Callister et al. (2021) to fit data of GWTC-3 for comparison. Specifically, the mean and width of the $\chi_{\rm eff}$ distribution change linearly with $q$ (referred to as the Base model); see Appendix B.2 for the detailed formula. The rate evolution model is the MD model (Madau & Dickinson, 2014), as defined in the Appendix B.3. We use both the non-/semi-parametric formula PowerLaw+Spline (PS; Edelman et al., 2022; Abbott et al., 2023a) and the popular parametric formula PowerLaw+Peak (PP; Talbot & Thrane, 2018; Abbott et al., 2021b) to model the primary-mass distribution, see Appendix B.1 for the details of mass functions. With the PP model, we find an anti-correlation between $\chi_{\rm eff}$ and $q$ distribution (i.e., $a<0$) at 99.2% credible level. However, the credibility decreases to 93% with the more flexible PS model, while the overall tendency remains unchanged, as shown in Figure 1.
@@
$\pi_{1}$ $(m_{1},\chi_{\rm eff}|{\bf\Lambda})=P_{1,m_{1}}(m_{1}|{\bf\Lambda})P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ (2)
@@
where $r_{2}$ is the mixed fraction of the second subpopulation, and $\mathcal{G}_{[-1,1]}$ is the truncated Gaussian distribution. $P_{m_{1}}$ and $\mathcal{PL}$ are the primary-mass functions for the two subpopulation, and we adopt $\mathcal{PS}$ and $\mathcal{PP}$ for $P_{m_{1}}$, respectively. $P(m_{2}|m_{1};{\bf\Lambda})$ is the normalized secondary-mass distribution conditioned on $m_{1}$, $P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ is the Base model that encodes the $\chi_{\rm eff}-q$ correlation (Callister et al., 2021), see Appendix B for the detailed formulas of the models. Note that we do not apply either the low-mass tapering function or the perturbation function to the power-law mass function for the second subpopulation, to avoid overly complex (bloated) models. Actually, with currently available data, the results are rather similar when the low-mass tapering and perturbation functions are applied, see Appendix C.
@@
$\pi_{\rm tran}$ $(m_{1},m_{2},\chi_{\rm eff}|{\bf\Lambda})=P_{m_{1}}(m_{1}|{\bf\Lambda})P(m_{2}|m_{1};{\bf\Lambda})$ (4)
@@ 3 Results
In Table 1, we summarize the Bayes factors of the novel models in this work compared to the Base (PS) and Base (PP) models, respectively. For both PS and PP cases, the Transition model and Mixture model are favored than the Base models. Furthermore, the Transition model and Mixture model without the $\chi_{\rm eff}$-$q$ correlation (i.e., $a=0,b=0$) are even more favored, with $\ln\mathcal{B}>4.2$. Figure 1(right panels) shows the posteriors of the hyperparameters describing the $\chi_{\rm eff}$ and $q$ correlation. In each case, the parameters of the Transition model and Mixture model are highly consistent with each other. For the PS case, the slope parameters $a$ and $b$ are consistent with zero. For the PP case, the $\chi_{\rm eff}$-$q$ correlation still remains, but the credibility of $a<0$ decreases to $\sim 90\%$ comparing to the Base (PP) model. It is consistent in both cases that the anti-correlation reduces when we introduce a second subpopulation of $\chi_{\rm eff}$ distribution, which means that the second subpopulation has contributed to the $\chi_{\rm eff}$-$q$ correlation in the whole population. Specifically, the anti-correlation of the entire BBH population mainly results from the superposition of two subpopulations. However, whether there is $\chi_{\rm eff}$-$q$ correlation in the first subpopulation remains uncertain. Additionally, we find there is no evidence for the $\chi_{\rm eff}$-$q$ correlation in the second subpopulation with currently available data, see Appendix D.2. In subsequent analysis, we adopt the results inferred with $a=0,b=0$, since the posterior distributions of parameters are broadly consistent with the results inferred with variable $a$ and $b$, see Appendix C.
@@
Base(PS) with $a=0$,$b=0$ 0
Transition(PS) 4.9
Transition(PS) with $a=0$,$b=0$ 6.7
Mixture(PS) 3.8
Mixture(PS) with $a=0$,$b=0$ 6.1
@@
$\ln{\mathcal{B}}_{\rm PP}^{X}$
Base(PP) 0
Base(PP) with $a=0$,$b=0$1.2
Transition(PP) 3.2
Transition(PP) with $a=0$,$b=0$ 4.9
Mixture(PP) 2.7
Mixture(PP) with $a=0$,$b=0$ 4.2
Note: All the log Bayes factors in the top / bottom are
relative to the Base (PS) / (PP) model that encode the
$\chi_{\rm eff}$-$q$ correlation and accompanied with the PS / PP
@@
Figure 1: Left&Mid: Constraints on the mean $\mu_{\chi}(q)$ and standard deviation $\sigma_{\chi}(q)$ of the $\chi_{\rm eff}$ distribution, as a function of BBH mass ratio $q$. The solid curves are the medians and the colored bands are the 90% credible intervals. Right: Posteriors of the hyperparameters describing the $\chi_{\rm eff}-q$ correlation. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
We find that the second subpopulation shows a significantly different $\chi_{\rm eff}$ distribution compared to the first subpopulation, see Figure 2. The divergence point of the Transition model in the primary-mass function is $m_{\rm t}=49^{+13}_{-9}M_{\odot}$  [^1] In this work, when quoting $X^{+Y}_{-Z}$, X means the posterior median and $[X-Y,Y+Z]$ means the central 90% credible interval., with a transition scale of $\delta_{\rm t}=6^{+3}_{-4}M_{\odot}$ ($\delta_{\rm t}=6^{+4}_{-4}M_{\odot}$) for the PS (PP) case. Note that $m_{\rm t}$ and $\delta_{\rm t}$ are degenerated with each other, and $\delta_{\rm t}$ is only loosely constrained, see Figure 3.
@@
As for the Mixture model, though the maximum mass of the first subpopulation is not well constrained, the masses of 99.5th and 99.75th, and 99.9th percentiles are better measured. Specifically, the $m_{99.9\%}$ of the first subpopulation is significantly different from that of the entire population, see Appendix C. The minimum mass of the second subpopulation is $\sim 30M_{\odot}$. These results are consistent with our previous analysis using spin magnitudes of BBHs (Wang et al., 2022; Li et al., 2024b), see Appendix C for the distributions of other related parameters. The $\chi_{\rm eff}$ distributions of both subpopulations are not symmetric with respect to zero, indicating the presence of formation channels other than star clusters.
@@
We find that $\mu_{\chi,2}>0$ at 98% (95%) credible level, and a symmetric $\chi_{\rm eff}$ distribution with respect to zero for the second population (i.e.,$\mu_{\chi,2}=0$) is disfavored by a Bayes factor of $\ln\mathcal{B}=1.4$ ($\ln\mathcal{B}=1.5$) for the PS (PP) case, which indicates that the hierarchical mergers cannot be produced (solely) by star clusters. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\rm eff,2}=0.40^{+0.32}_{-0.30}$ ($\mu_{\rm eff,2}=0.39^{+0.39}_{-0.37}$) for the PS (PP) case, which is consistent with the hierarchical mergers in the AGN disks (Yang et al., 2019). However, some other formation channels, like chemically homogeneous evolution (Mandel & de Mink, 2016) and stable mass transfer (Bavera et al., 2021) can not be ruled out based solely on the effective spin distribution (Zevin et al., 2021). We have also inferred using two more flexible model for the second $\chi_{\rm eff}$ distribution, one with variable edges for the truncated Gaussian, the other is spline function (Golomb & Talbot, 2023), see Appendix D.1. We find that the second $\chi_{\rm eff}$ distribution is not symmetric about zero, but having preference for positive values. Additionally, the negative values for the second $\chi_{\rm eff}$ distributions can not be ruled out.
@@
Figure 2: Effective-spin distributions of the two subpopulation inferred by the Transition and Mixture models with $a=0$,$b=0$. The lines represent the mean values and 90% credible intervals.
@@ 4 Conclusions and Discussion
In this work, we investigate the origins of $\chi_{\rm eff}$-$q$ correlation in the BBHs (Callister et al., 2021; Abbott et al., 2023a) with data of GWTC-3, using the dedicated models that introduce a second $\chi_{\rm eff}$ distribution for the second (high-spin) subpopulation.
In both PS and PP cases, the $\chi_{\rm eff}$-$q$ correlation significantly reduces when we include a second subpopulation for the $\chi_{\rm eff}$ distribution. Additionally, there is no evidence for a $\chi_{\rm eff}-q$ correlation in the second subpopulation, see Appendix D.2. Therefore we can conclude that the $\chi_{\rm eff}$-$q$ anti-correlation of the entire BBH population mainly results from the superposition of two subpopulations. However, whether there is $\chi_{\rm eff}$-$q$ correlation in the first subpopulation remains uncertain. Because for the PS case, $a$ is consistent with zero, while for the PP case, we still find $a<0$ at $\sim 90\%$ credible level, see Figure 1.
@@
The stable mass transfer formation channel is expected to produce BBHs with $\chi_{\rm eff}-q$ anti-correlation (Olejak et al., 2024; Banerjee & Olejak, 2024), while BBHs from the common envelop channel may exhibit an opposite correlation (Bavera et al., 2020, 2021). Therefore, it is possible that there is a $\chi_{\rm eff}$-$q$ correlation in the first subpopulation, which is consistent with the first-generation BBHs. Note that the first subpopulation may consist not only the isolated formation channels but also the dynamical formation channels (Ray et al., 2024; Li et al., 2024a). When the detections are enriched, more subpopulations may be recognized and the correlations of their parameters may also be measured.
The $\chi_{\rm eff}$ distribution of the second subpopulation (shown in Figure 2), is consistent with the hierarchical mergers in gas-rich environments, such as AGN disks (Yang et al., 2019). Some other formation channels may also produce $\chi_{\rm eff}$ distributions similar to that of the second subpopulation inferred with our models (see Zevin et al., 2021), such as chemically homogeneous evolution (Mandel & de Mink, 2016) and stable mass transfer (Bavera et al., 2021). However, the second subpopulation found in this work is actually consistent with the population of hierarchical mergers found in our previous works (Li et al., 2024b), where we have provided smoking-gun evidence, specifically a spin-magnitude distribution of approximately 0.7 (Gerosa & Fishbach, 2021). Therefore, hereafter we simply interpolate the second subpopulation as hierarchical mergers. Antonini et al. (2025) attributed this subpopulation to hierarchical mergers in star clusters, as they found that the $\chi_{\rm eff}$ distribution is consistent with a uniform distribution in the range of approximately (-0.5, 0.5). To check out whether such difference is caused by the systematic bias in our analysis, we perform end-to-end mock data studies including the parameter estimation and hierarchical analysis of mock signals, see Appendix E. We find that the two subpopulations in the $\chi_{\rm eff}$ distribution can be successfully recognized. Additionally, we repeat the simulation for many times, the recoveries of the second $\chi_{\rm eff}$ distributions are always broadly consistent with the injections. When injected distribution is symmetric (asymmetric), then the recovered distributions are almost symmetric (asymmetric). Thus, the identified asymmetry in the second subpopulation’s $\chi_{\rm eff}$ distribution in our study is attributed to the data itself rather than to methodological biases. With currently available data, we find the symmetric $\chi_{\rm eff}$ distribution of the second subpopulation ($\mu_{\chi,2}$) is less favored by $\ln\mathcal{B}\sim 1.5$. A more conclusive evidence may be found when the data from the fourth observing run [^2] See https://gracedb.ligo.org/superevents/public/O4/ for the catalog of detection candidates. are released.
Many simulations (e.g., Yang et al., 2019; Santini et al., 2023; Cook et al., 2024) show that the hierarchical mergers in disk-like environments can produce a $\chi_{\rm eff}$ distribution peaking at $\sim 0.4$ that is consistent with our results. Additionally, AGN-driven mergers (including first generations) will exhibit a $\chi_{\rm eff}$-$q$ correlation that is found in the entire BBH population (Santini et al., 2023; Cook et al., 2024). Therefore, it is very likely that AGN-disk channel has contributed to the BBHs detected by LVKC (Abbott et al., 2019, 2021a, 2024, 2023b). Some works also suggest that the BBHs with primary masses at $\sim 35M_{\odot}$ peak are consistent with dynamical formation channels in star clusters (Li et al., 2024a; Ray et al., 2024), as predicted by Antonini et al. (2023). Therefore, the star clusters have also contributed to the hierarchical mergers, if the remnants of the previous mergers are sufficiently retained in the environments (Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2024). Our further analysis with a more flexible models also show that the negative values for the second $\chi_{\rm eff}$ distributions can not be ruled out, see Appendix D.1.
@@
However, there are only $\sim 11$ hierarchical mergers (high-spin events) in GWTC-3 (Li et al., 2024b), making it difficult to determine the mixture fractions of AGN-disk like and star-cluster formation channels. We will address such issue when the GW data are significantly enriched. The fourth observing run (O4) of the LIGO-Virgo-KAGRA GW detectors is currently underway, and the number of detections is rapidly increasing (see https://gracedb.ligo.org/latest/). At the end of O4, more than four times as many events are expected to be observed compared to O3 (Callister, 2024). With the enriched data more subpopulations / formation channels of BBHs may be identified (Zevin et al., 2021), and the mixture fractions and the parameter correlations of subpopulations can be better determined (Li et al., 2022, 2024a; Guo et al., 2024; Heinzel et al., 2024; Callister, 2024).